Pi-6 Please don a rectangular boam. Wo-I.2KLF including beam Self weig . fl-poor ps 81 T 30 5
Answer #1

12x12=1.44626 IPUL = 6x1.6 9.6K tv t v 2-30 d = 30-3=27 8/ L 8/ Mu = 9-6X16+9.6811:44466= 153.60 +76.86184:32=414,72 K-H bal